An automotive manufacturer wants to know the proportion of new car buyers who prefer foreign cars over domestic.

Step 1 of 2:

Suppose a sample of 3190 new car buyers is drawn. Of those sampled, 2552 preferred domestic rather than foreign cars. Using the data, estimate the proportion of new car buyers who prefer foreign cars. Enter your answer as a fraction or a decimal number rounded to three decimal places.
Answer #1

Given that, a sample of 3190 new car buyers is drawn. Of those sampled, 2552 preferred domestic rather than foreign cars.

And number of samples who prefered foreign cars = 3190 - 2552 = 638

Therefore, the proportion of new car buyers who prefer foreign cars is 638/3190 = 0.200